OpenStudy (anonymous):

is |(-1,-2)| = (|-1|,|-2|)

OpenStudy (anonymous):

is that a point on a graph

OpenStudy (anonymous):

on properties of absolute value are or not?

OpenStudy (anonymous):

oops are they equal or not?

OpenStudy (anonymous):

I ment is (-1,-2) a point on a graph or an equation because if its a point on a graph then you take the absolute value of both numbers so yeah basically they would be equal (1 , 2)

OpenStudy (anonymous):

(1 , 2) would be the actual point

OpenStudy (anonymous):

let h(x)=(-sin(x),-cos(x)) then is |h(x)|=|(-sin(x),-cos(x))|=(sin(x),cos(x))?

OpenStudy (asnaseer):

I would imagine that |(-1, -2)| means "what is the magnitude of the vector (-1, -2)

OpenStudy (asnaseer):

are you studying vectors?

OpenStudy (anonymous):

yeah this question is very vague

OpenStudy (anonymous):

no ,is real analysis

OpenStudy (asnaseer):

the notation used still looks like it is asking for the magnitude of a vector. so, if h(x) = (-sin(x), -cos(x)) then:\[|h(x)|=|(-\sin(x), -\cos(x))|=\sqrt{\sin^2(x)+\cos^2(x)}=1\]

OpenStudy (anonymous):

actually i am trying to solve this

OpenStudy (anonymous):

wait a minute

OpenStudy (anonymous):

\[s={(x,y) \epsilon R ^{2}:x ^{2}+y ^{2}}\] S is the unit circle in the plane define \[h:[0,2\pi)\rightarrow S\] by \[h(\theta)=(\cos \theta , \sin )\]

OpenStudy (anonymous):

@jacobian you meant \(S=\{(x,y)\in\mathbb{R}^2:x^2+y^2=1\}\) anyways yes that is a valid coordinate map

OpenStudy (anonymous):

i've try it like this Let ϵ>0 and a∈R be given. Choose δ=ϵ. Then we have that \[|\sin \theta - \sin a|\le 2|\frac{ \theta - a }{ 2 }|\\ = |\theta - a| < \]

OpenStudy (anonymous):

this is also similar for cosine they are in the form of \[|\theta - a |<epsilon \]

OpenStudy (anonymous):

so if we choose \[\delta = \min \left\{ \delta _{1},\delta _{2}\ \right\} \] we get \[\left| h(\theta) - h(a)\right| = \left| (\cos \theta,\sin \theta) - (\cos a , \sin a)\right|\] then i don't know how to continue

OpenStudy (anonymous):

@asnaseer @amistre64 @KingGeorge

OpenStudy (anonymous):

are you trying to prove this function \(h:\mathbb{R}\to\mathbb{R}^2\) is continuous...?

OpenStudy (anonymous):

nop ,h:[0,2pi)->S